Nam Kading (01.01.2014 to 03.31.2014)

You are here

PDF versionPDF version
January - March
Communications activities: 

No communication activities

Unusual Events at the TEAM Site: 

No unusual events to report

New Species: 

No new species to report

Protocol Activities: 

Vegetation protocols

A total of 1148 stems were measured in 2 sample plots. This concludes the 2013/4 sampling period and all data has been entered onto the TEAM website.

Terrestrial vertebrate protocols

17 camera traps's were deployed (and collected) at one site, concluding the TV survey for 2013/4. Species identification is currently being verified before entry onto the TEAM website.

Climate station

The station's software was updated as per instructions, along with the installation of a new temperatire sensor.

Protocol Problems: 

No problems experienced

Schedule Problems: 

No scheduling problems.

Logistical Problems: 

No logistical problems


No suggestions at this stage.